# Anthropocentric bias and the possibility of artificial cognition

### v1

@@ 3.2 Type-II anthropocentrism
The key point is that competence, once abstracted from a narrowly human-centric understanding of the term, should be related to the generality and flexibility of the system’s computations rather than superficial resemblance to the human cognitive architecture. An LLM that relies on a giant lookup table of memorized addition problems would not be competent at addition, because it could not flexibly apply its memorized content-specific transitions to novel cases. But a neural network that learns a general addition algorithm could be seen as competent in the domain, even if the solution it has converged upon is distinctly unhuman-like.
@@ 3.3 An objection
We now anticipate an objection to our discussion of Type-II anthropocentrism: since the only indisputable examples of language-driven cognition we have are human, and since the capacities of LLMs are acquired through training on human-produced data, isn’t it appropriate to treat human cognition as the only appropriate yardstick for studying LLMs? Our response is that it depends on how we conceive of human cognition as a yardstick. While we can – and presumably must – begin our investigation of LLM competencies by comparing them to our own, the role of human cognition as a benchmark can be overstated in two ways. First, the human competencies we use as reference points should be regarded as what philosopher Ali Boyle calls “investigative kinds” (Boyle, 2024) - they serve as an initial search template, but not as necessary conditions for cognitive status. Second, questions of the form “Do LLMs have cognitive competence ?” should be treated as empirical questions. Though this may seem self-evident, it is a principle that is easy to violate, particularly when we assume that facts about implementation are among the distinguishing features between competence and competence . If facts about the physical implementation of a competence are included in its definition, then it will be impossible for LLMs, which are implemented in silicon computers, to acquire any of the competencies that humans enjoy. Moreover, this impossibility will be logical, rather than empirical. In order to preserve the empirical character of debates about LLM capacities, therefore, we must focus on the algorithmic level of description, where facts about implementation are explicitly set aside.
@@
The investigation of cognitive capacities in LLMs is best viewed as an iterative, cyclic process in which our conception of the relevant competencies and our understanding of the mechanisms that implement them in LLMs mutually inform and revise each other. Solving the challenge of mapping cognitive tasks to capacities often involves consideration of the underlying mechanisms (Francken et al., 2022). But to home in on the mechanisms responsible for a particular competence, we must make principled decisions about the level of abstraction at which to characterize the mechanism and about how to delineate the boundaries of the mechanism itself. These decisions, in turn, depend on how we conceptualize the cognitive phenomenon we are trying to explain. What results is a investigative process in which our characterization of cognitive tasks, our ontology of capacities, and our understanding of mechanisms evolve in tandem as we search for maximally predictive and explanatory mappings between them. In the case of LLMs, this process may lead us quite far from our initial starting point. Though we inevitably begin by searching for human-like competencies in LLMs, the gradual discovery and refinement of LLM-specific mechanisms may ultimately produce a novel ontology of cognitive kinds, one that is optimized for explaining the distinctive strengths and weaknesses of machine intelligence rather than human intelligence. In this way, the iterative nature of the investigative process allows it to drift away from its anthropocentric origins.
@@ 4 Conclusion
Like humans, LLMs can be right for the wrong reasons; we should not take good performance on various benchmarks, particularly those designed without attention to construct validity and potential confounds, at face value. However, they can also be wrong for the wrong reasons. Various auxiliary factors can interfere with their performance, such that both successes and failures only provide defeasible evidence about their underlying competence in a domain. Our analysis suggests that anthropocentrism, being more subtle than anthropomorphism, has garnered less theoretical scrutiny. Despite this relative neglect, we argued that anthropocentric bias can significantly impede the objective assessment of LLM capabilities. To address this gap, we presented a systematic taxonomy of anthropocentric reasoning about LLMs.
The most direct way to work out whether auxiliary factors actually cause performance degradation on a task is to look inside the model. As such, mechanistic interpretability techniques are essential for counteracting Type-I anthropocentrism. However, the role of mechanistic interpretability in addressing Type-II anthropocentrism is more nuanced. It is possible to decode a feature or circuit of interest from a large neural network even if that feature or circuit does not significantly contribute to the network’s functionality (Makelov et al., 2023; Huang et al., 2023). Consequently, there is a risk of anthropomorphic projection in mechanistic interpretability research. As previously noted, we should be cautious not to combat anthropocentrism by resorting to anthropomorphism. Nevertheless, there exist mechanisms that process information in ways that differ from typical human strategies but are nonetheless robust and general. Mechanistic interpretability research can help identify these mechanisms. When such mechanisms are discovered, we should not dismiss them simply because they deviate from common human approaches.

### v2

@@ 3.2 Mechanistic chauvinism
The key point is that competence, once abstracted from a narrowly human-centric understanding of the term, should be related to the generality and flexibility of the system’s computations rather than superficial resemblance to the human cognitive architecture. An LLM that relies on a giant lookup table of memorized addition problems would not be competent at addition, because it could not flexibly apply its memorized content-specific transitions to novel cases. But a neural network that learns a general addition algorithm could be seen as competent in the domain, even if the solution it has converged upon is distinctly non-human.
@@ 3.3 An objection
We now anticipate an objection to our discussion of mechanistic chauvinism: since the only indisputable examples of language-driven cognition we have are human, and since the capacities of LLMs are acquired through training on human-produced data, isn’t it appropriate to treat human cognition as the only appropriate yardstick for studying LLMs? Our response is that it depends on how the yardstick is used. While we can – and presumably must – begin our investigation of LLM competencies by comparing them to our own, the role of human cognition as a benchmark can be overstated in two ways. First, the human competencies we use as reference points should be regarded as what philosopher Ali Boyle calls “investigative kinds” (Boyle 2024) - they serve as an initial search template, but not as necessary conditions for cognitive status. Second, questions of the form “Do LLMs have cognitive competence ?” should be treated as empirical questions. Though this may seem self-evident, it is a principle that is easy to violate, particularly when we assume that facts about implementation are among the distinguishing features between competence and competence . If facts about the physical implementation of a competence are included in its definition, then it will be impossible for LLMs, which are implemented in silicon computers, to acquire any of the competencies that humans enjoy. Moreover, this impossibility will be logical, rather than empirical. In order to preserve the empirical character of debates about LLM capacities, therefore, we must focus on the algorithmic level of description, where facts about implementation are explicitly set aside.
@@
The investigation of cognitive capacities in LLMs is best viewed as an iterative, cyclic process in which our conception of the relevant competencies and our understanding of the mechanisms that implement them in LLMs mutually inform and revise each other. Solving the challenge of mapping cognitive tasks to capacities often involves consideration of the underlying mechanisms (Francken et al. 2022). But to home in on the mechanisms responsible for a particular competence, we must make principled decisions about the level of abstraction at which to characterize the mechanism and about how to delineate the boundaries of the mechanism itself. These decisions, in turn, depend on how we conceptualize the cognitive phenomenon we are trying to explain. What results is a investigative process in which our characterization of cognitive tasks, our ontology of capacities, and our understanding of mechanisms evolve in tandem as we search for maximally predictive and explanatory mappings between them. In the case of LLMs, this process may lead us quite far from our initial starting point. Though we inevitably begin by searching for human-like competencies in LLMs, the gradual discovery and refinement of LLM-specific mechanisms may ultimately produce a novel ontology of cognitive kinds, one that is optimized for explaining the distinctive strengths and weaknesses of machine intelligence rather than human intelligence. In this way, the iterative nature of the investigative process allows it to drift away from its anthropocentric origins.
@@ 4 Conclusion
Like humans, LLMs can be right for the wrong reasons; we should not take good performance on various benchmarks, particularly those designed without attention to construct validity and potential confounds, at face value. However, they can also be wrong for the wrong reasons. Auxiliary factors can suppress performance in ways that obscure latent competence, so both successes and failures offer only defeasible evidence. While anthropomorphism has received substantial critical attention, anthropocentric bias has gone largely unexamined. We have argued that it poses a serious obstacle to the objective assessment of LLM capacities, and we offered a taxonomy of its forms to help guide future research.
The most direct way to work out whether auxiliary factors actually cause performance degradation on a task is to look inside the model. As such, mechanistic interpretability techniques are essential for counteracting auxiliary oversight. However, the role of mechanistic interpretability in addressing mechanistic chauvinism is more nuanced. It is possible to decode a feature or circuit of interest from a large neural network even if that feature or circuit does not significantly contribute to the network’s functionality (Makelov et al. 2024, Huang et al. 2023). Consequently, there is a risk of anthropomorphic projection in mechanistic interpretability research. As noted above, we should not combat anthropocentrism by resorting to anthropomorphism. Nevertheless, there exist mechanisms that process information in ways that differ from typical human strategies but are nonetheless robust and general. Mechanistic interpretability research can help identify these mechanisms. When they are discovered, we should not dismiss them simply because they deviate from the human case. Doing so would foreclose the possibility of understanding artificial cognition on its own terms.